OpenStudy (anonymous):

write an equation in the point slope form of the line that passes through the given point and has the given slope m ? (3,5)m=1

OpenStudy (anonymous):

substitute the given values for x1 and y1....into the following form of a line.... y-y1=m(x-x1).....where x1=3 and y1=5

OpenStudy (anonymous):

then substitute in m=1 for the m....and you are done.